Example 3.8: Two electric Furnaces are supplied with single-phase circuit at 100 V from

& 3-phase, 13,750 V supply by means of two single-phase Scott-connected transformers with

similar seco;z dary windings when the load on main transformer is 625 kW and on the teaser

is 1000 kW, zohat current will flow in each of 3-phase lines at unity power factor ?


1
Expert's answer
2020-11-13T10:52:07-0500

Transformer coefficient "k=\frac{N_1}{N_2}=\frac{13750}{100}=137.5"

Then "\frac{\sqrt{3}}{2}\times 137.5=119.08"

The furnace current is "I_a=\frac{625\times 10^3}{100\times 1}=6250" A

Similarly, "I_b=\frac{1000\times 10^3}{100\times 1}=10000" A

On the three-phase side, "I_A=\frac{6250}{119.08}=52.49" A.